# A Coincidence Algebra Bundle for Decay Quivers: An Algebraic Approach to Gamma-ray Spectroscopy

### v1

@@ Abstract
Motivated by the need for a more comprehensive algebraic structure to calculate coincidence probabilities of a general decay scheme for gamma ray spectroscopy, we model the decay scheme, rather naturally, as a quiver through which we define a decay quiver. The path algebra of quivers is the underlying, more general, algebra for transition matrices that is typically used in modeling decay schemes. The path algebra allows for concatenation of transitions which affords the calculation of cascade probabilities. We extend the path algebra to allow for the multiplication of non-composable paths, i.e., transition that don’t directly share a level connecting them. We define the coincidence algebra as the algebra that allows for such an extension and realize it as the fibres for a coincidence algebra bundle, the base space of which is the path algebra where decay schemes live. A given decay schemes coincidence probabilities are calculated on its fibre. Detection maps are defined as linear maps on the base space that map transition probabilities to detected probabilities.
keywords:
Gamma-ray Spectroscopy , Fibre Bundles , Coincidence Summing , Path Algebra
[…]
[label2]organization=Department of Physics, addressline= University of Guelph, city=Guelph, postcode=N1G 2W1, country=ON, Canada
1 Introduction
@@
An algebraic structure for spectroscopic analysis is proposed that models the gamma decay as a quiver with a defined structure; a quiver is a directed graph upon which a vector space and algebra is composed [Assem_Skowronski_Simson_2006]. It is common practice to treat nuclear levels using a graph theoretical approach. However, the manner in which that is done is through transition matrices/ adjacency matrices [SEMKOW1990437], which take a matrix representation of the more fundamental algebra built from a quiver; the path algebra [Assem_Skowronski_Simson_2006]. The path algebra constructed from a quiver that models a decay scheme allows for probability calculations for coincidence events (as do the transition matrices, though the path algebra keeps cascade terms separate and does not sum with the direct transitions). The conventional path algebra, however, is restricted to consecutive paths; that is, directly connected paths. In this work, the path algebra is extended into a tensor algebra where non-consecutive products may be calculated. This serves as a more general algebraic structure for nuclear level schemes, which may assist in the calculation of coincidence probabilities for gamma-ray spectroscopy.
1.1 Transition Matrices
Conventional treatment for decay diagrams is done with transition matrices [SEMKOW1990437]. General decays can be modeled using strictly lower triangular matrices, where the matrix algebra allows for the calculation of cascade probabilities via powers of the transition matrix.
@@ 1.1 Transition Matrices
For a general decay scheme, as shown in Figure 1, the transition probabilities can be stored in the transition matrix
$x=\begin{pmatrix}0&\\ x_{10}&0\\ x_{20}&x_{21}&0\\ \vdots&\vdots&\vdots\\ x_{n0}&x_{n1}&x_{n2}&\cdots&x_{nn-1}&0\end{pmatrix}$, (1)
and the branching ratios are stored in a branching vector
$f=(f_{0},f_{1},f_{2},\cdots,f_{n}).$ (2)
Taking powers of the transition matrix will give cascade probabilities on the order of the power in the matrix position of the direct transition. For example, all two gamma cascade probabilities can be calculated from:
$x^{2}=\begin{pmatrix}0&\\ 0&0\\ x_{10}x_{21}&0&0\\ x_{10}x_{31}+x_{32}x_{20}&x_{21}x_{32}&0&0\end{pmatrix}$. (3)
To collect all orders of cascade probabilities, from single transitions to $n$ gamma cascades, we may power expand the transition matrix:
$\overline{x}^{n}=\sum_{k=0}^{n}x^{k}.$ (4)
With the branching vector determining the level at which a given decay starts, the probability that a transition from level $k$ to level $l$ will occur for a given decay can be given as follows:
$P_{TM}(x_{kl})=\left[diag\left(f\overline{x}^{n}\right)x\right]_{kl}.$ (5)
The diagonal matrix $diag\left(f\overline{x}^{n}\right)$ gives the probability that a given level is reached. To calculate the probability that two transitions, $k\rightarrow l$ and $q\rightarrow m$, occur we take the probability of the first transition to occur multiplied by the probability that it leads to the second with
$P_{TM}(x_{kl}\cap x_{qm})=\left[diag\left(\left[f\overline{x}^{n}\right]_{i}\right)x\right]_{kl}\left[diag\left(\left[\overline{x}^{n}\right]_{li}\right)x\right]_{qm}.$ (6)
@@
The diagonal matrix $diag\left(\left[\overline{x}^{n}\right]_{li}\right)$ gives the probability that from level $l$ the decay leads to a given level below it. The above probabilities are probabilities of emission, simply that the transition occurs. In gamma-ray spectroscopy we must deal with the probability of detection. It is the map between detected probabilities and emitted probabilities that is the practice of gamma-ray spectroscopy. To model this, the transition matrices can be mapped to include different detection probabilities, such as efficiencies. This is done in [SEMKOW1990437] where observed summing probabilities are given; this paper is motivated by [SEMKOW1990437] where we look to further generalize the mathematical procedure of coincidence probabilities. The work of Semkow et.al. [SEMKOW1990437] first established the transition matrix formalism for nuclear decay schemes. In their paper, their motivation was to calculate the true coincidence summing corrections that occur in gamma-ray spectroscopy. When detecting the emission of gamma rays in a decay, there is some non zero probability that two gammas from a cascade can hit your detector making it appear like the detection of a single gamma-ray with energy being the sum of the two. If the direct transition of that cascade exist, then extra counts are added to the peak of the direct transition and counts are lost from the the two smaller energy gamma peaks. When this occurs, we may refer to the loss of counts due to summing as summing out and the gain of counts as summing in. It is crucial in gamma-ray spectroscopy that the coincidence summing effect is well know in order to arrive at an accurate measurement. The transition matrix formalism was extended to include different phenomena such as coincidence summing with X-rays and $511$ keV gamma-rays from positron annihilation [KORUN1993478, RICHARDSON1990344]. The work of [JUTIER20071344] takes a unique approach to modeling coincidence summing with hints of a graph theory, but falls short of reaching the rigorous formalism of quivers and path algebra. Recent work by Schmidt [schmidt2025measurabilitytruecoincidencesumming] gives a more comprehensive history of coincidence summing and also extends the matrix formalism of [SEMKOW1990437] for 180 degree coincidence ( as a correction for coincidence summing) and for gated (coincidence) spectra. The work presented here is a further generalization of the transition matrix formalism for coincidence probabilities in nuclear level schemes.
@@
The transition matrices are limited to creating product probabilities between connected transitions. This is due to the algebra of strictly lower triangular matrices. As shown in [SEMKOW1990437], this concatenation of probabilities allows for coincidence summing calculations of connected transitions. To calculate the coincidence summing of non-connected transitions, one must extend the algebra. It is the coincidence summing of non-connected transition in particular that motivates this work. To tackle this problem, we propose a structure that offers to deal with coincidence probabilities in the most general sense. The transition matrices emerge out of a more general algebra, called the path algebra of quivers. The transition matrices are isomorphic to the path algebra modded by the ideal of relations that equate all cascades to their direct transition. We extend this path algebra to a restricted tensor path algebra we call the coincidence algebra.
2 Quivers and Path Algebra
@@
$(a|\alpha_{1},...\alpha_{l}|b)\cdot(c|\beta_{1},...\beta_{k}|d)=\delta_{bc}(a|\alpha_{1},...\alpha_{l},\beta_{1},...\beta_{k}|b).$ (9)
@@
$\mathbb{1}_{\mathcal{Q}}=\sum_{a\in\mathcal{Q}_{0}}\widehat{\epsilon}_{a}.$ (10)
@@
where $K\mathcal{Q}_{l}$ is the subspace whose basis $\mathcal{P}_{l}$ is the set of all paths of length $l$. The arrow ideal $R_{\mathcal{D}}$ , which is defined as isomorphic to the quotient vector space $K\mathcal{D}/K\mathcal{D}_{0}$, may be decomposed into source subspaces $\mathcal{S}_{i}$ or target subspaces $\mathcal{S}_{i}$ composed of basis vectors with their source or target as level $i$, respectively:
@@
2.1 Decay Quiver
Definition 2.10.
@@ 2.1 Decay Quiver
The bijection between $\mathcal{D}_{0}$ and $\{0,1,2,...n-1\}$ gives the number of arrows in the quiver, that is, the cardinality of the arrow set, to be $|\mathcal{D})|=n(n-1)/2$. We denote the path  xacxcdxnb
@@
$P_{PA}(\widehat{p}_{i})=\left\langle V_{t}\left(b\cdot\overline{\tau}^{n}\right)\cdot d,\widehat{p_{i}}\right\rangle_{p},$ (26)
@@
$P_{PA}(\widehat{p}_{i}\cap\widehat{p}_{j})=\left\langle V_{t}\left(b\cdot\overline{\tau}^{n}\right)\cdot\tau,\widehat{p_{i}}\right\rangle_{p}\left\langle V_{t}\left(\epsilon_{t(\widehat{p}_{i})}\cdot\overline{\tau}^{n}\right)\cdot\tau,\widehat{p}_{j}\right\rangle_{p},$ (27)
@@
where $\tau=d-b$ defined in equation 24, that is, the decay vector with the branching terms subtracted off. In this formulation, the target vertex projector acts as the diagonalization (left multiplying) of the decay vector with the branching from above terms multiplied through. To select a given transition the path equivalence bilinear form is used. It should be noted, given that the target vertex projector gives back a vector of stationary paths, when multiplied onto the decay vector $d$, what is given is a vector in $k\mathcal{D}_{1}$ still. Therefore, when find a given path with the path form only single terms are produced– no equivalence between different path lengths is produced yet, as would be given by the path form in general. To arrive at notions of coincidence summing, which in this formalism we may generally state as the path equivalence between paths of different lengths, one must power expand the decay vector which $V_{t}$ is being multiplied onto to create the high order path terms. Algebraically, one may think of coincidence summing as path equivalence; however, a more complete depiction of coincidence summing is more aptly posed in the context of detection maps.
2.2 Detection Maps
@@
The decay vectors describing nuclear levels schemes live in $k\mathcal{D}$, or more accurately $k\mathcal{D}_{0}\oplus k\mathcal{D}_{1}\subset k\mathcal{D}$ such that their decay norm is 1. The decay vectors have components that are the probabilities of emission– they are intrinsic to the decay nuclei. However, there exist vectors in this space that do not satisfy the unit vector condition but whose components can still be seen as probabilities. From this, we may consider probabilities related to detection also living in this space. When we measure the decay of a nucleus, of course what we actually detect from our spectra are not the complete and correct transition probabilities but a different set of probabilities that have been mapped from the transition probabilities via the experiment or manner of detection. One may consider the probability of detecting a transition with a detector efficiency, and a vector that contains the efficiencies can also be taken to exist in the path algebra. If we compute probabilities via equation 26 with a detection mapped decay vector, where efficiencies are now multiplied onto the transition probabilities, what we get is not the probability of emission but the probability of detecting that gamma/transition. In general, for more realistic observations of decay nuclei, there are other factors beside detector efficiency that go into the detected probabilities that one doing gamma spectroscopy must consider when attempting to map their way back to the transition probabilities. What we posit here is that detected probabilities merely still live in the path algebra, and we can consider a number of different detection maps that take the decay vectors describing nuclear decay schemes into different vectors that account for different probabilities involved in detection. Since for any two vectors, we can always find a linear map from one to the other ( either with some simple construction or if needed a change of basis) we may consider detections maps in general to be maps $\varphi\in End(k\mathcal{D})$. For this present work we do not worry about the exact form of a given map – say for full energy detection using peak efficiencies, detection with total efficiencies, or for probability of not being detected at any energy [^2] These are maps [SEMKOW1990437] considers for the transition matrices to calculate coincidence summing corrections. – and merely speak of the detection maps by there operation. With this, we may now give a path algebraic formulation for the probability vector that provides the probability of detection for a given transition where summing in and summing out effects are considered via detections maps $\varphi_{i}$ and $\varphi_{o}$, respectively: [^3] This is analogous to the matrix $S=NAM$ given in [SEMKOW1990437].
$\Gamma(d)=V_{t}\left(b\cdot\overline{\varphi_{o}(\tau)}^{n}\right)\cdot\left[\overline{\varphi_{i}(\tau)}^{n}-\mathbb{1}\right]\cdot V_{s}\left(\overline{\varphi_{o}(\tau)}^{n}\right).$ (28)
@@
In contrast from equation 26, the target vertex projector now contains probabilities multiplied onto vertices(stationary paths) that account for summing out from above, $\overline{\varphi_{i}(d)}^{n}$ contains the paths with summing in given via the power expansion, and the source vertex projector now gives summing out from below. All of which, when given together as shown above, provide a vector where the probability of detecting a transition is given by the component of the corresponding basis vector/path. However, different from the transition matrix formalism where the summing in effects are added with the direct transition, the summing in terms exist in separate paths, i.e. they have their own basis vector. To add them back together, the path equivalence form, which equates paths with the same source and target – but does not distinguish between path length, can be used to calculate probabilities for detecting a given transition with summing effects accounted for:
@@
Up to here, this new formulation just restates the transition matrix formalism with the caveat that summing in terms are kept separate. The path equivalence quotient path algebra, that is, the path algebra with the path equivalence modded out, is isomorphic to the transition matrices. However, this new formalism may now be extended beyond the path concatenation that we are restricted to with the transition matrices and also allow for a greater separability between equivalent paths. If we wish to allow for not only paths – being connected arrows – but pairs, or even sets, of non-connected arrows, we must extend the algebra to a tensor decay space defined in the following section
3 Coincidence Algebra Bundle
3.1 Coincidence Tensor Space

### v2

# thanks: A footnote to the article title

@@ Abstract
Motivated by the need for a more comprehensive algebraic structure to calculate coincidence probabilities of a general decay scheme for gamma ray spectroscopy, we model the decay scheme, rather naturally, as a quiver through which we define a decay quiver. The path algebra of quivers is the underlying, more general, algebra for transition matrices that is typically used in modeling decay schemes. The path algebra allows for concatenation of transitions which affords the calculation of cascade probabilities. We extend the path algebra to allow for the multiplication of non-composable paths, i.e., transition that don’t directly share a level connecting them. We define the coincidence algebra as the algebra that allows for such an extension and realize it as the fibres for a coincidence algebra bundle, the base space of which is the path algebra where decay schemes live. A given decay schemes coincidence probabilities are calculated on its fibre. Detection maps are defined as maps on the base space that map transition probabilities to detection probabilities.
preprint: APS/123-QED
I Introduction
@@
An algebraic structure for spectroscopic analysis is proposed that models the gamma decay as a quiver with a defined structure; a quiver is a directed graph upon which a vector space and algebra is composed [2]. It is common practice to treat nuclear levels using a graph theoretical approach. However, the manner in which that is done is through transition matrices [8, 5], which take a matrix representation of the more fundamental algebra built from a quiver–the path algebra [2]. The path algebra constructed from a quiver that models a decay scheme allows for probability calculations for coincidence events (as do the transition matrices, though the path algebra keeps cascade terms separate and does not sum with the direct transitions). The conventional path algebra, however, is restricted to consecutive paths; that is, directly connected paths. In this work, the path algebra is extended into a tensor algebra where non-consecutive products may be calculated. This serves as a more general algebraic structure for nuclear level schemes, which may assist in the calculation of coincidence probabilities for gamma-ray spectroscopy.
I.1 Transition Matrices
Conventional treatment for decay diagrams is done with transition matrices [8]. General decays can be modeled using strictly lower triangular matrices, where the matrix algebra allows for the calculation of cascade probabilities via powers of the transition matrix.
@@ I.1 Transition Matrices
The diagonal matrix $diag\left(\left[\overline{x}^{n}\right]_{li}\right)$ gives the probability that from level $l$ the decay leads to a given level below it. The above probabilities are probabilities of emission; they indicate that the transition occurs. In gamma-ray spectroscopy we must deal with the probability of detection. It is the map between detected probabilities and emitted probabilities that is the practice of gamma-ray spectroscopy. To model this, the transition matrices can be mapped to include different detection probabilities, such as efficiencies. This is done in [8] where observed summing probabilities are given; this paper is motivated by [8] where we look to further generalize the mathematical procedure of coincidence probabilities. The work of Semkow et.al. [8] first established the transition matrix formalism for nuclear decay schemes, though largely building off the work of [1, 6]. In their paper, their motivation was to calculate the true coincidence summing corrections that occur in gamma-ray spectroscopy. When detecting the emission of gamma rays in a decay, there is some non zero probability that two gammas from a cascade can hit your detector making it appear like the detection of a single gamma-ray with energy being the sum of the two. If the direct transition of that cascade exist, then extra counts are added to the peak of the direct transition and counts are lost from the the two smaller energy gamma peaks. When this occurs, we may refer to the loss of counts due to summing as summing out and the gain of counts as summing in. It is crucial in gamma-ray spectroscopy that the coincidence summing effect is well know in order to arrive at an accurate measurement. The transition matrix formalism was extended to include different phenomena such as coincidence summing with X-rays [5]. The work of [4] takes a unique approach to modeling coincidence summing with hints of a graph theory, but falls short of reaching the rigorous formalism of quivers and path algebra. Recent work by Schmidt [7] gives a more comprehensive history of coincidence summing and also extends the matrix formalism of [8] for 180 degree coincidence ( as a correction for coincidence summing) and for gated (coincidence) spectra. The work presented here is a further generalization of the transition matrix formalism for coincidence probabilities in nuclear level schemes.
@@
The transition matrices are limited to creating product probabilities between connected transitions. This is due to the algebra of strictly lower triangular matrices. As shown in [8], this concatenation of probabilities allows for coincidence summing calculations of connected transitions. To calculate the coincidence summing of non-connected transitions, one must extend the algebra. It is the coincidence summing of non-connected transition in particular that motivates this work. To tackle this problem, we propose a structure that offers to deal with coincidence probabilities in the most general sense. The transition matrices emerge out of a more general algebra, called the path algebra of quivers. The transition matrices are isomorphic to the path algebra modded by the ideal of relations that equate all cascades to their direct transition. We extend this path algebra to a restricted tensor path algebra we call the coincidence algebra.
II Quivers and Path Algebra
@@
$(a|\alpha_{1},...\alpha_{l}|b)\cdot(c|\beta_{1},...\beta_{k}|d)=\delta_{bc}(a|\alpha_{1},...\alpha_{l},\beta_{1},...\beta_{k}|d).$ (3)
@@
$\mathbb{I}_{\mathcal{Q}}=\sum_{a\in\mathcal{Q}_{0}}\widehat{\epsilon}_{a}.$ (4)
@@
where $K\mathcal{Q}_{l}$ is the subspace whose basis $\mathcal{P}_{l}$ is the set of all paths of length $l$. The arrow ideal $R_{\mathcal{D}}$ , which is defined as isomorphic to the quotient vector space $K\mathcal{D}/K\mathcal{D}_{0}$, may be decomposed into source subspaces $\mathcal{S}_{i}$ or target subspaces $\mathcal{T}_{i}$ composed of basis vectors with their source or target as level $i$, respectively:
@@
II.1 Decay Quiver
Definition II.10.
@@ II.1 Decay Quiver
The bijection between $\mathcal{D}_{0}$ and $\{0,1,2,...n-1\}$ gives the number of arrows in the quiver, that is, the cardinality of the arrow set, to be $|\mathcal{D}|=n(n-1)/2$. We denote the path  xacxcdxnb
@@
$P(\widehat{p}_{i})=\left\langle V_{t}\left(b\cdot\overline{\tau}^{n}\right)\cdot d,\widehat{p_{i}}\right\rangle_{p},$ (20)
@@
$P(\widehat{p}_{i}\cap\widehat{p}_{j})=\left\langle V_{t}\left(b\cdot\overline{\tau}^{n}\right)\cdot\tau,\widehat{p_{i}}\right\rangle_{p}\langle V_{t}(\epsilon_{t(\widehat{p}_{i})}\cdot\overline{\tau}^{n})\cdot\tau,\widehat{p}_{j}\rangle_{p},$ (21)
@@
where $\tau=d-b$ as defined in equation 24, that is, the decay vector with the branching terms subtracted off. In this formulation, the target vertex projector acts as the diagonalization (left multiplying) of the decay vector with the branching from above terms multiplied through. To select a given transition the path equivalence bilinear form is used. It should be noted, given that the target vertex projector gives back a vector of stationary paths, when multiplied onto the decay vector $d$, what is given is a vector in $k\mathcal{D}_{1}$ still. Therefore, when finding a given path with the path form only single terms are produced– no equivalence between different path lengths is produced yet, as would be given by the path form in general. To arrive at notions of coincidence summing, which in this formalism we may generally state as the path equivalence between paths of different lengths, one must power expand the decay vector which $V_{t}$ is being multiplied onto to create the high order path terms. Algebraically, one may think of coincidence summing as path equivalence; however, a more complete depiction of coincidence summing is more aptly posed in the context of detection maps.
II.2 Detection Maps
@@
For a given decay n-quiver, a decay vector living in the path algebra models a nuclear level scheme with n levels. The decay vectors components give the probability for the transition corresponding to basis vector path. As such, the decay vectors are units vectors under the decay norm. However, there of course exist vectors in the path algebra that do not correspond to transition probabilities satisfying the norm condition, but may still be considered as probabilities. In this manner, detection vectors may also live in the path algebra where the components not only include the transition probabilities but also efficiencies or different detection probabilities. We posit then that the path algebra for a given quiver is the space off all detection nuclear level schemes, such that transition probabilities are mapped to detection probabilities in this space. If we compute probabilities via equation 26 with a detection mapped decay vector, where efficiencies are now multiplied onto the transition probabilities, what we get is not the probability of emission but the probability of detecting that gamma/transition. In general, for more realistic observations of decaying nuclei, there are other factors beside detector efficiency that go into the detected probabilities that one doing gamma spectroscopy must consider when attempting to map their way back to the transition probabilities; one can consider a number of different detection maps that take the decay vectors describing nuclear decay schemes into different vectors that account for different probabilities involved in detection. A detection map that corresponds to the full energy hit of a gamma may be given by,
$\varphi_{h}(\tau)=\sum_{\widehat{p}\in\mathcal{P}_{1}}e^{p}_{\widehat{p}}\langle\widehat{p},\tau\rangle_{p}\widehat{p}$, (22)
where $e^{p}_{\widehat{p}}$ is the peak efficiency for the gamma associated to the path $\widehat{p}$ of length one. We may also consider the avoidance of a gamma at any energy ( full detection or Compton scattered). To do so, however, we must consider the detector geometry. When talking about gammas summing together and avoiding summing out one must consider angular correlations: the probability that two gammas are emitted at a certain angle. For now, we just consider isotropic emission such that the probability for two gammas hitting the same detector is just given by the number of detectors. Therefore, we may map a decay vector to a detection vector that accounts for the probability of it hitting a given detector at any energy:
$\varphi_{a}(\tau)=\sum_{\widehat{p}\in\mathcal{P}_{1}}\frac{e^{t}_{\widehat{p}}}{N}\langle\widehat{p},\tau\rangle_{p}\widehat{p}$, (23)
for $N$ detector ( one may imagine N detectors spherically placed around an emission source) where $e^{t}_{\widehat{p}}$ is the total efficiency. To arrive at a notion of summing out, we consider the probability that a given gamma does not hit a given detector at any energy:
$\varphi_{o}(\tau)=\sum_{\widehat{p}\in\mathcal{P}_{1}}\left[1-\frac{e^{t}_{\widehat{p}}}{N}\right]\langle\widehat{p},\tau\rangle_{p}\widehat{p}$, (24)
These are all maps analogous to those defined in [8] with the transition matrix formalism and used in the literature afterwards to discuss summing corrections. To account for the summing in, cascade probabilities were built via the power expansion of the decay vector. However, if the power expansion creates the gammas of interest with cascade terms to account for summing in, we must now consider the detector geometry involved for the summing in. Maintaining our isotropic, summing in probabilities can just be considered with the number of detectors ([8] considers only one detector so the efficiencies are sufficient in accounting for summing probabilities. For multiple detectors one must extend this of course). We then redefine the power expansion used for detection mapped decay vectors with isotropic emission.
$\overline{\varphi(\tau)}^{n}_{N}=\sum_{k=0}^{n}\frac{\tau^{k}}{N^{k+1}},$ (25)
with $k$ the degree of summing. We may now give a path algebraic formulation for the probability vector that provides the probability of detection for a given transition where summing in and summing out effects are considered: [^1] This is analogous to the matrix $S=NAM$ given in [8] with the rate $R$ taken out.
$\Gamma(d)=V_{t}\left(b\cdot\overline{\varphi_{o}(\tau)}^{n}\right)\cdot\left[\overline{\varphi_{h}(\tau)}^{n}_{N}-\mathbb{I}_{Q}\right]\cdot V_{s}\left(\overline{\varphi_{o}(\tau)}^{n}\right).$ (26)
@@
In contrast from equation 26, the target vertex projector now contains probabilities multiplied onto vertices (stationary paths) that account for summing out from above, $\overline{\varphi_{h}(\tau)}^{n}_{N}$ contains the paths with summing in given via the power expansion, and the source vertex projector now gives summing out from below. All of which, when given together as shown above, provide a vector where the probability of detecting a transition is given by the component of the corresponding basis vector/path. However, different from the transition matrix formalism where the summing in effects are added with the direct transition, the summing in terms exist in separate paths, i.e. they have their own basis vector. To add them back together, the path equivalence form, which equates paths with the same source and target – but does not distinguish between path length, can be used to calculate probabilities for detecting a given transition with summing effects accounted for:
@@
Up to here, this new formulation just restates the transition matrix formalism with the caveat that summing in terms are kept separate; the path equivalence quotient path algebra, that is, the path algebra with the path equivalence modded out, is isomorphic to the transition matrices. However, this new formalism may now be extended beyond the path concatenation that we are restricted to with the transition matrices and also allow for a greater separability between equivalent paths. If we wish to allow for not only paths – being connected arrows – but pairs, or even sets, of non-connected arrows (i.e., coincidences in general) in the algebra (as shown in Figure 2), we must extend the algebra to a tensor decay space defined in the following section
Figure 2: General decay diagram showing the connected transitions A and B and the non-connected or coincidence transitions A and C.
III Coincidence Algebra Bundle
III.1 Coincidence Tensor Space